# Re-visiting Skip-Gram Negative Sampling: Dimension Regularization for More Efficient Dissimilarity Preservation in Graph Embeddings

### v1

@@ 5.2. Non-contrastive Self-Supervised Learning
In this work, we provide a new perspective on dissimilarity preservation in graph representation learning and show that dissimilarity preservation can be achieved via dimension regularization. Our main theoretical finding shows that when node repulsion is most needed and embedding dot products are all increasing, the difference between the original skip-gram dissimilarity loss and a regularizer that induces origin-centered dimensions vanishes. Combined with the efficiency of dimension operations over node repulsions, dimension regularization bypasses the need for SGNS. We then introduce a generic algorithm augmentation that prioritizes positive updates, given that real-world graph data often contain missing edges (Young et al., 2021), and when node repulsion is needed, utilizes dimension regularization instead of SGNS. Our experimental results show that the augmented versions of LINE and node2vec preserve the link-prediction performance of the original algorithms while reducing runtime by over $60$%. In fact, for several real-world graphs, removing dissimilarity preservation altogether performs well; however for synthetic graphs with high connectivity, our augmentation clearly outperforms baselines.
@@ Appendix A Proofs
where the 4th line follows from the cyclic property of the trace.
A.2. Proposition 2.3
@@
(25) $X_{i}^{t+1}$ $\approx X_{i}^{t}+\eta\sum_{j\in N(i)}\left(1-\frac{1}{2}\left((X_{i}^{t})^{T}X_{j}^{t}\right)\right)X_{j}^{t}$
@@
(26) $P^{\prime}_{SG}=\|\mathbf{1}_{S>0}\odot\left(\vec{1}\vec{1}^{T}-\frac{1}{2}XX^{T}\right)\|_{F}^{2}$
From Gunasekar et al. (2017), gradient descent for matrix completion is implicitly regularized to yield the minimum nuclear norm (lowest rank) stationary point. In the case where $G$ is connected, minimizing the nuclear norm implies that gradient descent on (26) causes $XX^{T}$ to approach $2\vec{1}\vec{1}^{T}$, and thus the dot products between all pairs of embeddings increases and $\mathcal{C}\to 2$.
Phase 2: Asymptotic. For constriction values $c\leq 2$, the proof is complete after Phase 1. On the other hand, for constriction values $c>2$, we can complete the proof by showing that following alignment, characterized as positive dot products for all pairs of embeddings, constriction monotonically increases with each gradient descent epoch. We show the monotonic increase in embeddings via induction. For any pair of embeddings $X_{i}^{t}$ and $X_{j}^{t}$, the dot product after a single gradient descent epoch is:
(27) $X_{i}^{t+1},X_{j}^{t+1}\right>=&\left(X_{i}^{t}+\eta\sum_{k\in N(i)}\sigma\left((X_{i}^{t})^{T}X_{k}^{t}\right)X_{k}^{t}\right)^{T}\\ &\left(X_{j}^{t}+\eta\sum_{k\in N(j)}\sigma\left((X_{j}^{t})^{T}X_{k}^{t}\right)X_{k}^{t}\right)\end{split}$
At time step $t^{\prime}$, the constriction $\mathcal{C}>0$, thus, once all of the dot products are expanded, we have $X_{i}^{t+1},X_{j}^{t+1}\right>=\left<X_{i}^{t},X_{j}^{t}\right>+C$, where $C>0$. Thus the dot product between any pair of embeddings strictly increases at time step $t^{\prime}+1$. By definition, the constriction remains positive, so by induction, we have shown that for all $t>t^{\prime}$, the constriction is monotonically increasing. ∎
@@ A.2.1. Supplemental Figures
Figure 7. High-level overview of proof for Proposition 2.3, which guarantees embedding collapse when only attraction updates are applied. In the beginning the embeddings are initialized near the origin, all with norm at most $b$. Then, in Phase 1, attraction update rule is approximately gradient descent for matrix completion; given that the latter is implicitly regularized to yield low-rank solutions the embeddings converge in direction. In the second phase, the embeddings asymptotically distance away from the origin in the same direction.
@@
Figure 7 provides a high-level summary of the proof for Proposition 2.3. In the beginning, the embeddings are initialized with norm $\leq b\leq\sqrt{2}$. Then, the embeddings converge in direction given that near the origin, the gradient of $P_{SG}$ is the gradient of a matrix completion problem. Further, gradient descent for matrix completion near the origin is implicitly regularized to yield the lowest rank solution, hence a convergence in the embedding direction. Thereafter, once the dot products between all pairs of nodes are positive in the “Asymptotic” phase, $\mathcal{C}$ becomes monotonically increasing.
We also empirically validate Proposition 2.3 in Figure 8. We initialize Erdös-Rényi graphs ($n=100$) of various densities, randomly initialize the embeddings around the origin, and then apply the positive-only update rule in (23). Figure 8 shows that even when the edge density is $0.05$, the constriction eventually becomes monotonically increasing.
@@
(30) $\|\nabla N^{\prime}_{SG}-N_{SG}\|^{2}_{F}$ $=\sum_{i=1}^{n}\Big{\|}\sum_{j^{\prime}\in\{j|S_{ij}>0\}}\sigma\left(X_{i}^{T}X_{j^{\prime}}\right)X_{j^{\prime}}\Big{\|}^{2}_{2}$
@@
Where in the above, $\beta_{\text{max}}$ is a constant and the upper bound on embedding norm squared and $m$ is the number of non-zero entries in $S$.
@@
Combining the bounds on the numerator and denominator together, we have:
@@
Where the inequality follows reducing the value of the denominator. This concludes the proof. ∎
@@
The gradient for $N^{\prime}_{SG}$, defined in Equation (10), is:
@@
The constriction $\mathcal{C}$ is the minimum value of the matrix K. From lemma A.1, we know that as constriction increases, the difference between $1$ and each of the entries of $K$ vanishes exponentially. Thus, there is a vanishing difference between $\frac{1}{2}\nabla_{X}N^{\prime}_{SG}$ and $\mathbf{1}X$, where $\mathbf{1}$ is the $n\times n$ all-ones matrix. Note that $\mathbf{1}X$ is also the gradient of the dimension regularizer $R$ introduced in (13). Putting these together we have:
@@
Thus, as constriction $\mathcal{C}$ increases, the difference between the gradient of $\frac{1}{2}N^{\prime}_{SG}$ and $R$ vanishes exponentially. By extension, the difference between gradient descent on $N^{\prime}_{SG}$ with a step-size of $\eta$ and gradient descent on $R$ with a step-size of $2\eta$ vanishes with increasing $\mathcal{C}$. ∎
@@ Appendix B Experimental Methodology
B.1. Graph Generation and Pre-processing
For all of the graph datasets listed in Table 2, we discarded all components except for the largest connected component. For the SBM graphs used in Figure 6, the number of nodes was fixed at $1000$, the number of blocks was fixed at $10$, and the in-group edge probability $p$ was fixed at $0.2$.
B.2. Runtime Evaluation
We perform batched gradient descent for both LINE II and node2vec II. In both cases, we construct batches of positive pairs where each batch consists of $1000$ pairs. The number of batches is $5n$, where $n$ is the number of nodes in the graph.
To determine the runtime of vanilla node2vec, we utilized the Grover and Leskovec (2016) code to generate the random walks but then performed the positive and negative updates in Python instead of using gensim which is implemented in C. This created a level playing field to compare against our augmentation which is implemented purely in Python. To simulate the random walks, we used the simulate_walks function in the Grover and Leskovec (2016) code, where we sampled 10 random walks of length 80 for each node in the graph. It is important to note that for all of the precision and transductive precision vanilla node2vec results, we utilized the gensim embeddings generated by the Grover and Leskovec (2016) code. We confirmed that the node2vec embeddings we generated for the runtime evaluation yielded similar precision results as the gensim embeddings.
All of our experiments were written in Python and executed on machines with Intel Xeon E5-2690 CPUS, 2.60 GHz, 30 MB of cache.
B.3. Initialization
For almost all algorithms, we uniformly randomly initialize the embeddings $X^{\text{init}}$ in $\left(0,1/(d)\right)^{n\times d}$. The only exception is for the Positive Only baselines, which are most stable when we initialize the embeddings $X^{\text{init}}$ uniformly in $\left(-1/(2d),1/(2d)\right)^{n\times d}$, but ultimately still collapse.
B.4. Hyperparameters
[…]
Table 3. Optimal hyperparameters
Graph Original Algorithm $\lambda$ $n_{\text{negative}}$
LastFM node2vec 100 1
LastFM LINE 100 1
Facebook node2vec 100 1
Facebook LINE 100 1
ca-HepTh node2vec 100 1
ca-HepTh LINE 100 1
Protein-Protein node2vec 100 1
Protein-Protein LINE 100 1
Wikipedia node2vec 1 100
Wikipedia LINE 1 100
We note that if the regularization update is performed too often (small $n_{\text{negative}}$) and $\lambda$ is large, the embeddings can diverge; specifically, instead of re-centering around the origin, the embeddings overshoot the origin and “reflect” about the origin, growing in magnitude each time. In our experiments, the embeddings produced numerical overflows when $\lambda=100$ and $n_{\text{negative}}\leq 10$ for the Facebook and Protein-Protein graphs; the Wikipedia graph embeddings overflowed when $n_{\text{negative}}=1$ and $\lambda=100$.
B.5. Transductive Precision
To calculate the transductive precision metric, we begin by splitting the original graph edge set into training and test edge sets, following a 70/30 split. The training graph is then the largest connected component induced by the training edge set. The embeddings are trained on the training graph and then evaluated on the test graph, which has the same node set as the training graph. The transductive precision@k value is then the average transductive top-k precision among all nodes with at least one neighbor in the test graph; for nodes with degree less than $k$ in the test graph, we set $k=d$ , where $d$ is the test degree of the node. When calculating the precision for a single node, all edges present in the training set are excluded from the top-k selection process.
@@ Appendix C Supplemental Evaluation
Runtime vs. Performance
Figure 5 only shows the augmented algorithm’s results using the optimal hyperparameter configurations, so in Figure 9 we include the runtime results for all hyperparameter configurations, as denoted by the orange Xs. The runtime for our augmentation decreases as $n_{\text{negative}}$ increases as the regularization update is performed less frequently; when $n_{\text{negative}}=1000$, the runtime is similar to the runtime for the positive-only baselines as nearly only positive updates are provided. For the LastFM, Facebook, ca-HepTh, and Protein-Protein graphs, the optimal hyperparameter configurations set $n_{\text{negative}}=1$ so the runtime for the optimal configuration is higher than that for most configurations while remaining less than the vanilla algorithm runtimes. Only for the Wikipedia graph does the optimal configuration yield a runtime comparable to the positive-only baselines.
To supplement the transductive precision results, Figure 10 characterizes all algorithms based on average link-prediction AUC-ROC. The AUC-ROC for a single node is calculated by taking the test-set edges as ground truth labels ($1$ for neighbors and 00 for non-neighbors) and the dot products between embeddings as prediction scores. The results in Figure 10 confirm the transductive precision results and show that our augmentation reduces runtime while preserving downstream performance.
Figure 9. To supplement Figure 3, this figure shows the runtime and transductive precision for all hyperparameter configurations of our augmentation algorithm. The optimal configuration is still denoted by the orange star but all other configurations are denoted by orange Xs. The non-optimal hyperparameter configurations yield faster training times but also worse performance.
Figure 10. The above figure shows the downstream performance of each algorithm as measured by average link-prediction AUC-ROC as opposed to transductive precision. To calculate the AUC-ROC for each node, we take edges in the test set as positive examples and non-edges as negative examples; the prediction scores are the dot products between pairs of embeddings. The AUC-ROC results confirm that our augmentation preserves out-of-sample performance while reducing runtime.
Performance vs. Node Features
SGNS traditionally samples nodes with probability proportional to $d^{3/4}$ where $d$ is the degree of the node. Because our dimension regularization augmentation defaults to treating nodes uniformly, it is important to assess whether there are correlations between node features and downstream performance. For instance, do low-degree nodes perform better under our augmentation compared to SGNS? Figures 11 and 12 plot downstream performance as a function of node degree, triangle count, and clustering coefficient. Overall, the relationship between node features and downstream performance is quite similar across the algorithms; the main exception, as noted previously, is that for the Wikipedia graph, our augmentation as well as the normalized positive-only baseline perform best. The lack of differences among algorithms is surprising and suggests that for these networks sampling negative samples proportional to $d^{3/4}$ performs similarly as sampling uniformly.
Figure 11. Because the SGNS node sampling probability is a function of node degree and our augmentation samples uniformly by default, it is important to assess whether the algorithms differ in their correlations between node features and downstream performance. For each node feature (degree, triangle count, and clustering coefficient), nodes are ranked and divided into $50$ bins, and then the average transductive precision is plotted for each bin. Overall, we see that the algorithms share similar correlations between node features and downstream performance despite differences in the node-sampling distributions.
Figure 12. The setup for the above figure is analogous with Figure 11 but for node2vec. The takeaways for node2vec are largely the same as those for LINE: the algorithms share similar correlations between node features and downstream performance.
Re-weighting Nodes
As discussed in subsection 2.2.1, our augmentation is also capable of a flexible weighting of nodes when performing repulsion just as SGNS offers this flexibility. While all other empirical evaluations of our augmentation weights nodes uniformly (using (13) over (15)), in Figure 13 we apply a weight of $d^{3/4}$ to each node as done in traditional SGNS. Figure 13 shows that while our augmentation offers flexible re-weighting, the alternative weighting yields results that are very similar to the uniform weighting results shown in the paper.
(a) Uniform Weighting
(b) Weight Nodes by $\text{degree}^{3/4}$
Figure 13. While our augmentation is capable of weighting nodes when performing repulsion, just as SGNS weights nodes as a function of degree, applying the traditional SGNS node weights (weight proportional to $\text{degree}^{3/4}$) via (15) does not significantly impact downstream performance compared to using uniform weights, which we use in all other empirical evaluations. In the figure, the curves for LINE II in the top subfigure correspond to setting $\vec{p}=\vec{1}$ whereas the LINE II curves in the bottom subfigure correspond to setting the weight of each node to $d^{3/4}$, where $d$ is the node’s degree. The curves for all other variations of LINE are identical in the top and bottom subfigures.

### v2

@@ 7. Conclusion
We provide a new perspective on dissimilarity preservation in graph representation learning and show that dissimilarity preservation can be achieved via dimension regularization. Our main theoretical finding shows that when node repulsion is most needed and embedding dot products are all increasing, the difference between the original skip-gram dissimilarity loss and the dimension-mean regularizer vanishes. Combined with the efficiency of dimension operations over node repulsions, dimension regularization bypasses the need for SGNS. We then introduce a generic algorithm augmentation that prioritizes positive updates. When node repulsion is needed and collapse approaches, the augmentation utilizes dimension regularization instead of SGNS. Our experimental results show that replacing SGNS with dimension regularization in LINE and node2vec preserves the link-prediction performance of the original algorithms while reducing runtime by up to $23.4$% and memory by up to $33.3\%$ for OGB benchmark datasets. Furthermore, for LINE, removing dissimilarity preservation altogether, a special case of our augmentation framework, reduces training time by $70.9\%$ on average while also increasing link prediction performance–especially for graphs that are globally sparse but locally dense. However, for graphs with low local density, repulsion is needed, and dimension regularization is an efficient and effective approach.
@@ Appendix A Proofs
Eq. (20) follows from the cyclic property of the trace.
A.2. Proof for Proposition 3.3
@@
(25) $X_{i}^{t+1}$ $\approx X_{i}^{t}+\frac{\eta}{2}\sum_{j\in N(i)}\left(1-\frac{1}{2}\left((X_{i}^{t})^{T}X_{j}^{t}\right)\right)X_{j}^{t}$
@@
(26) $P^{\prime}_{SG}=\|\mathbf{1}_{S>0}\odot\left(\mathbf{1}-\frac{1}{2}XX^{T}\right)\|_{F}^{2}$
From Gunasekar et al. (2017), gradient descent for matrix completion is implicitly regularized to yield the minimum nuclear norm (lowest rank) stationary point. In the case where $G$ is connected, minimizing the nuclear norm implies that gradient descent in Eq. (26) causes $XX^{T}$ to approach $2\mathbf{1}$, and thus the dot products between all pairs of embeddings increase and $\mathcal{C}\to 2$.
Phase 2: Asymptotic. To complete the proof we show that once $\mathcal{C}>0$, the constriction monotonically increases with each gradient-descent update to $P_{SG}$. For any pair of embeddings $X_{i}^{t}$ and $X_{j}^{t}$, the dot product after a single gradient descent epoch is:
(27) $X_{i}^{t+1},X_{j}^{t+1}\right>=&\left(X_{i}^{t}+\eta\sum_{k\in N(i)}\sigma\left(-(X_{i}^{t})^{T}X_{k}^{t}\right)X_{k}^{t}\right)^{T}\\ &\left(X_{j}^{t}+\eta\sum_{k\in N(j)}\sigma\left(-(X_{j}^{t})^{T}X_{k}^{t}\right)X_{k}^{t}\right)\end{split}$
If $\mathcal{C}>0$ at $t$, then all of the dot-product terms from distributing the right-hand side of Eq. (27) are positive, and we have $X_{i}^{t+1},X_{j}^{t+1}\right>=\left<X_{i}^{t},X_{j}^{t}\right>+\delta$, where $\delta>0$. Thus, the dot product between any pair of embeddings strictly increases from step $t$ to $t+1$, and therefore the constriction monotonically increases if $\mathcal{C}>0$ at step $t$. ∎
@@ A.2.1. Supplemental Figures
[…]
[Illustration of proof for Proposition 3.3]The figure shows the geometry of graph embeddings as characterized in the proof. The embeddings are shown for three stages: initialization, alignment, and asymptotic. At initialization, the embeddings are shown as a cluster of black points near the origin of an x-y axis. At alignment, the points have clustered at a point on the circumference of a circle centered at the origin with radius equal to the square-root of 2. At the asymptotic stage, the points are clustered at a point even further from the origin.
Figure 5. High-level overview of proof for Proposition 3.3, which guarantees embedding collapse when only attraction updates are applied. In the beginning, the embeddings are initialized near the origin. Then, in Phase 1, the attraction update rule is approximately gradient descent for matrix completion; given that the latter is implicitly regularized to yield low-rank solutions the embeddings converge in direction. In the second phase, the embeddings asymptotically distance away from the origin in the same direction.
[…]
[Validation of Proposition 3.3]A line plot where the x-axis is the gradient descent step and the y-axis is the Constriction. Each line represents a different Erdos-Renyi graph density. For each density p (0.5, 0.2, 0.1, 0.05), the line starts at the origin and then eventually becomes monotonically increasing with the step count. The curves where p is 0.5, 0.2, or 0.1 increase sharply during the initial gradient descent steps before a more gradual increase. For the smallest p (0.05), the curve initially dips before monotonically increasing above zero.
@@
Figure 5 provides a high-level summary of the proof for Proposition 3.3. In the beginning, the embeddings are initialized with norm $\leq b\leq\sqrt{2}$. Then, the embeddings converge in direction given that near the origin, the gradient of $P_{SG}$ is approximately the gradient of a matrix completion problem. Further, gradient descent for matrix completion near the origin is implicitly regularized to yield the lowest rank solution, hence a convergence in the embedding direction. Thereafter, once the dot products between all pairs of nodes are positive in the “Asymptotic” phase, $\mathcal{C}$ becomes monotonically increasing.
We also empirically validate Proposition 3.3 in Figure 6. We initialize Erdös-Rényi graphs ($n=100$) of various densities, randomly initialize the embeddings around the origin, and then apply the positive-only update rule in Eq. (23). Figure 6 shows that even when the edge density is $0.05$, the constriction eventually becomes monotonically increasing.
@@
(30) $\|\nabla N^{\prime}_{SG}-\nabla N_{SG}\|^{2}_{F}$ $=\sum_{i=1}^{n}\Big{\|}\sum_{j^{\prime}\in\{j|S_{ij}>0\}}\sigma\left(X_{i}^{T}X_{j^{\prime}}\right)X_{j^{\prime}}\Big{\|}^{2}_{2}$
@@
Where in the above, $\beta_{\text{max}}$ is a constant and the upper bound on embedding norm squared, and $m$ is the number of non-zero entries in $S$.
@@
Combining the bounds on the numerator and denominator, we have:
@@
Where the inequality follows from reducing the value of the denominator. This concludes the proof. ∎
@@
The gradient for $N^{\prime}_{SG}$, defined in Eq. (10), is:
@@
The constriction $\mathcal{C}$ is the minimum value of the matrix K. From lemma A.1, we know that as constriction increases, the difference between $1$ and each of the entries of $K$ vanishes exponentially. Thus, there is a vanishing difference between $\frac{1}{2}\nabla_{X}N^{\prime}_{SG}$ and $\mathbf{1}X$, where $\mathbf{1}$ is the $n\times n$ all-ones matrix. Note that $\mathbf{1}X$ is also the gradient of the dimension regularizer $R$ introduced in Eq. (13). Putting these together:
@@
Thus, as $\mathcal{C}$ increases, the difference between the gradient of $\frac{1}{2}N^{\prime}_{SG}$ and $R$ vanishes exponentially. By extension, the difference between gradient descent on $N^{\prime}_{SG}$ with a step-size of $\eta$ and gradient descent on $R$ with a step-size of $2\eta$ vanishes with increasing $\mathcal{C}$. ∎
@@ Appendix B Experimental Methodology
Hyperparameter Optimization.
To reduce the search space, we sequentially optimize the learning rate, node2vec random-walk, and augmentation parameters following the routine below. The optimal values are documented in our code repository.
  1. (1)
    Learning rate. For each dataset and vanilla algorithm (LINE and node2vec), select the learning rate that maximizes the validation set AUC-ROC among $\eta\in\{0.001$, $0.01$, $0.1$, 1.0}1.0\}1.0 }.
  2. (2)
    node2vec random-walk parameters. For each dataset, run a node2vec grid search over $p,q\in\{0.25,0.50,1.0,2.0,4.0\}$. For Cora, CiteSeer, and PubMed, select the values $p,q$ that maximize the validation AUC-ROC. For the OGB graphs, select the values that maximize the metric corresponding to the OGB leaderboard (see Table 5).
  3. (3)
    Augmentation parameters. For each graph and vanilla model, perform a grid search over $n_{\text{negative}}\in\{5,10,100\}$ and $\lambda\in\{0.1,1.0,10.0,100.0\}$. Select the values that maximize the same metrics optimized in the previous step. For Cora, CiteSeer, and PubMed, select the top-five highest-scoring parameter values; for ogbl-collab, select the top four; for the remaining OGB graphs, select the top three. In the final test-set evaluation, the results for II correspond to the best-performing values of $n_{\text{negative}}$ and $\lambda$, as measured on the test set using the metrics specified in step (2).
@@ Appendix C Supplemental Evaluation
OGB metrics.
In addition to the AUC-ROC metrics shown in the main body, in Table 5, for the OGB graphs, we also include the metrics reported on the OGB leaderboards.
Table 5. Performance for OGB graphs according to leaderboard metrics.
Graph Metric Model Performance
I II0 II
ogbl-collab Hits@50 node2vec 0.25 0.14 0.17
LINE 0.05 0.14 0.19
ogbl-ppa Hits@100 node2vec 0.15 0.02 0.02
LINE 0.01 0.04 0.08
ogbl-citation2 MRR node2vec 0.32 0.07 0.16
LINE 0.03 0.03 0.08
ogbl-vessel AUC-ROC node2vec 0.96 0.83 0.87
LINE 0.50 0.50 0.50
Non-uniform negative sampling.
Table 6. When SGNS and our augmentation are instantiated with non-uniform negatives, our augmentation still preserves link-prediction performance as measured by AUC-ROC. Column I is SGNS instantiated with the common configuration of $\alpha=3/4$ and $k=5$. Column II is our augmentation instantiated with $\vec{p}$ as the probability vector of $P_{3/4}$ (see Section 3.2.1). Entries marked as “-” denote out-of-memory errors on OGB datasets when $k=5$, highlighting the GPU memory cost of SGNS.
Graph node2vec AUC-ROC LINE AUC-ROC
I II I II
CiteSeer 0.68 0.79 0.61 0.62
Cora 0.84 0.85 0.60 0.66
PubMed 0.76 0.84 0.69 0.72
ogbl-collab - 0.96 0.71 0.78
ogbl-ppa - 0.98 0.92 0.76
ogbl-citation2 - 0.94 0.67 0.58
ogbl-vessel - 0.88 0.50 0.50
In Section 3.2.1, we note that SGNS introduces the two hyperparameters $\alpha$ and $k$. $\alpha$ enables non-uniform sampling over nodes, where the probability of sampling a node is proportional to the node degreeα. $k$ specifies the number of negative samples per positive pair. In the main body, we analyze both SGNS and our augmentation with the uniform distribution ($\alpha=0$, $k=1$, $\vec{p}=\vec{1}$) as this is a common instantiation of SGNS. However, in practice, it is also common to set $\alpha=3/4$ and $k=5$, based on promising empirical results from Mikolov et al. (2013). In Table 6, we show that when we instantiate SGNS and our augmentation with non-uniform sampling, our augmentation preserves link-prediction performance. When we set $k=5$ for SGNS, we obtain out-of-memory (OOM) errors on the OGB graphs, illustrating the memory cost of node contrast.